OpenStudy (anonymous):

16ax + 4x2 +16a2 factor please?

OpenStudy (anonymous):

16ax+4x2+16a2 = 4(x2+4ax+4a2) =4(x+2a)^2

OpenStudy (anonymous):

Gahhh i still don't understand.. is there like an easy way of explaining this to me??

OpenStudy (anonymous):

4x^2+16ax+16a^2= 4(x^2+4ax+4a^2) we want to factor what's in the parentheses ax^2+bx+c a*c=4 we need 2 factors of 4 than add to 4, those factors are 2 & 2 4(x^2+2ax+2ax+4a^2)= 4[x(x+2a)+2a(x+2a)]= 4[(x+2a)(x+2a)]= 4(x+2a)^2

OpenStudy (anonymous):

look up "factor by grouping" or the "ac method"

OpenStudy (anonymous):

Haha i was just reading the ac method.. Thanks! :) I'll see if i figured it out... and i'll message you if i need more help! :)

OpenStudy (anonymous):

you lost me right here we want to factor what's in the parentheses ax^2+bx+c a*c=4

OpenStudy (anonymous):

i just wrote ax^2+bx+c because thats the standard form of a quadratic

OpenStudy (anonymous):

you can have anything for a, b, or c, but its letting you know that a is the coefficient of x^2, b is the coefficient of x, and c is the constant

OpenStudy (anonymous):

alright thanks :)